# MECD+: Unlocking Event-Level Causal Graph Discovery for Video Reasoning

### v3

@@ Abstract
Index Terms:
Causal discovery, Causal reasoning, Video understanding, Video reasoning.
@@ I Introduction
Video causal reasoning aims to understand and analyze video content from a causal perspective. It requires models to comprehend temporal relationships, anticipate actions, and adapt to dynamic visual elements across frames. This capability is essential in various real-world applications, including autonomous driving [1], activity recognition [2], video surveillance [3], and even complex decision-making scenarios like robotic navigation [4]. Among these applications, Video Question Answering (VQA) [5, 6, 7, 8, 9] represents one of the most prominent tasks in video causal reasoning, where models are tested on their causal ability to understand videos through causal questions such as explanations, predictions, and counterfactual assumptions. Traditional VQA tasks can be viewed as attempts to discover single causal links within videos, as they typically identify one reason from given multiple options that explain the event described in the question.
Recent advancements in VQA have expanded beyond traditional tasks by incorporating more sophisticated causal reasoning methods. Notable trends include enhancing causal chain inference from simple one-to-one relations to multi-to-one chains, as seen in multiple correct answers VQA [10, 11, 12], and improving answer causal discovery through temporal grounding techniques, as seen in Grounded Video Question Answering (grounded VQA) [13, 14, 15]. Recent research also endeavors to infer a more comprehensive causal graph focusing on a particular event within its contexts [16, 17]. Moreover, REXTIME [18] takes an alternative approach by identifying causal relations between video events and generating targeted QA questions based on the causal relations, though it primarily focuses on a limited set of causal chains.
@@
Despite recent advancements, as shown in Tab. I, current video causal reasoning tasks remain limited in scope, primarily focusing on question-answering-based approaches that discover a single causal relation within a video. However, when the scene to be analyzed is more complex, these tasks often fail to meet the required understanding standards. For example, in a video where a person slips and falls while carrying a tray of drinks, the cause might be combined with a wet floor, slippery shoes, and losing balance due to walking too quickly. Besides, these tasks often lack the ability to perform fine-grained event-level reasoning, which is crucial in real-world scenarios. Such detailed reasoning typically leads to clearer insights into video content. Most importantly, they fail to provide a comprehensive, structured causal representation. For instance, in traffic surveillance videos, a detailed analysis of events occurring at different times is essential to determine which events, or combinations of events, resulted in a specific accident among the chain of traffic accidents.
@@
(1) Causality confounding indicates that the original causal relationships between events are disrupted or interfered with by other relay or adjacent events. Such confounding stems from the fact that many causal relationships flow through an intermediary event that acts as a bridge. As shown in Fig. 1 (c), event “submitting the paper” serves as a necessary bridge between “taking the test” and “obtaining a grade.” In this case, this bridge event might be mistakenly regarded as the only cause of the result event, while another cause, “taking the test,” is overlooked. However, the bridge event can only occur if “taking the test” happens first. Therefore, we cannot identify the real causality between events that linked by such bridges following a simple predictive criterion, and eliminating such confounding is thus crucial for an accurate causal discovery.
@@
(2) Illusory Causality, which includes illusory temporal and existence causality. Illusory temporal causality exists when events exhibit a close correlation in temporal distribution. Such correlation may mislead the test of real causality. As shown in Fig. 1 (d), the event “adding oil when cooking” often occurs before “adding vegetables to stir-fry,” but there is no real causality between them. As for illusory existence causality, it occurs when some objects in early events may serve as necessary existence conditions of a later event. For instance (Fig. 1 (e)), consider determining the causal relation between “a large brown dog enters the room” (at the start of the video) and “the dog runs towards the camera.” (at the end of the video). Although the presence of the dog in the former event is a prerequisite for the subsequent event, it does not directly cause the dog to rush towards the camera.
@@
Building upon the preceding discussion, we introduce a novel framework named VGCM (Video Granger Causality Model) to tackle MECD. This framework executes the Event Granger Test via an efficient mask-based event prediction model. It deduces the causality of a premise event by comparing the predicted features of the result event when the premise is either masked or unmasked. Furthermore, to mitigate the challenges of causality confounding and illusory causality discussed earlier, we integrate two additional causal inference techniques—front-door adjustment [22, 23, 24] and counterfactual inference [22, 25, 26]—into our framework. Specifically, these techniques compensate for or remove the causal effects of previous or subsequent adjacent bridge events to eliminate confounding. Simultaneously, they address the illusory causality issue by incorporating the chain of thoughts [27, 28, 29] and existence-only descriptions during inference.
Furthermore, We enhance the model’s robustness by utilizing context chain reasoning when conducting the Event Causality Test through further interaction between in-context causal chains. In reality, there are numerous instances where partial causes, when combined, contribute to the formation of the complete cause leading to the result. For example, high school student Sam’s math scores have been declining, despite his late-night study sessions. This causal relation becomes clearer when factoring in the context that his math class is the first period in the morning, as staying up late likely hinders his ability to focus during class, contributing to his lower scores. Therefore, we conduct the Event Causality Test by randomly masking several premise events simultaneously during the training phase.
@@
This manuscript is an extension of our NeurIPS 2024 paper [30], we make the following contributions:
  • The release and deeper analysis of MECD+, a larger multi-video source dataset.
  • An enhancement of the Event Causality Test through robust context chain reasoning and efficient non-regressive complete graph reasoning.
  • Additional metrics and further analysis of the MECD task, such as input modalities, hallucination issues, and downstream applications.
@@ II Related work
Video Causal Reasoning. Many existing tasks have tried to carry out causal reasoning in videos. Among these, the most common task is Video Question Answering (VQA), aiming to deduce a reason according to the question, grounded-QA methods such as SeViLA, NextGQA and Momentor [7, 13, 31] grounded a single reason happens before the result. However, the traditional VQA task does not extend to abducting multiple reasons, merely abducting a single causal chain from reason to the provided result. Recently, some methods, such as Glance [32], VideoTree [14], and TimeCraft [15] have constructed more detailed causal chains through reasoning from coarse-grained to fine-grained levels or bidirectional paths. Although these works improve the causal reasoning process, they only abduct the causal chain for a single result.
@@
Furthermore, many tasks are based on VQA for further causal reasoning attempts. Physical state reasoning task CLEVRER [9], CATER [33], and V-CDN [34] explored causal reasoning based on physics and other basic laws in virtual scenes. However, these tasks haven’t been committed to extending to the general video causal reasoning. Neural-symbolic paradigm AAR [35] and LMLN [36] symbolized data and derived inference rules using external knowledge. However, they can only reason within a defined symbol domain. The most similar task VAR [17] predicted explanation events with premise events, and the causality is introduced during its prediction process. However, it hasn’t been committed to discovering the complete causal graph. Another concurrent study REXTIME [18] determined the causal relation between two events in a video and then designed QA questions targeted at the relations. Although it explicitly used causal relationships, it only utilized a limited number of causal chains rather than inferring a complete causal graph.
@@
The task of causal reasoning in natural language processing (NLP) generally involves inferring the causal relations between words within a text paragraph. Methods usually incorporate external knowledge, either through knowledge graphs or prompts. For example, RFBFN [43] and CauSeRL [44] utilized knowledge related to candidate causal events and external causal statements from the knowledge graph. However, these approaches are generally limited to handling sentences with a single causal relation pair.
@@ III Multi-Event Causal Discovery
Figure 3: Statistics of the MECD Dataset. As shown in Figures (a1) and (a2), our dataset mainly analyzes videos that are longer than 120 seconds and contain five or more events. As shown in Figures (b1) and (b2), the proportion of causal and non-causal videos in the MECD is relatively balanced, moreover emphasizing the existence of causal relations between adjacent events.
To quantize the ability about causal discovery of a given model in multi-event videos, we propose the task Multi-Event Causal Discovery (MECD). In formulation, given a video $\mathcal{E}$ that contains chronologically organized $N$ events, $\mathbb{E}:=\{e_{1},\dots,e_{N}\}$, this task aims at determining whether any previous event $e_{n}$ ($n<N$) has a causal relation with the last one (i.e., $e_{N}$). Specifically, an event $e_{n}=\{v_{n},c_{n}\}$ consists of a video clip $v_{n}$ and the corresponding caption $c_{n}$. Without loss of generality, relations of previous events to the last one can be expressed as ${\bm{r}}=[r_{1},\dots,r_{N-1}]$, where $r_{k}$ ($k<N$) is set to “1” to indicate the existence of $e_{k}$’s causal relation with $e_{N}$, and “0” otherwise. Notably, this setting is generalizable to causal relations of any of two events as long as we cut off the video and treat the latter one as the last event.
@@ III-A Task Data
The ActivityNet Captions dataset [45] is built on ActivityNet v1.3 which includes 20k 120-second YouTube untrimmed videos. ActivityNet Captions [45] is widely used in dense video captioning tasks with the caption and timestamp of each sub-event. The EgoSchema dataset [46] contains over 5,000 very long-form video language understanding questions spanning over 250 hours of real, diverse, and high-quality egocentric video data. NExTVideo [8] consists of 6,000 videos that are longer and richer in objects and interactions in daily life. All these three datasets are extensively utilized in video question answering tasks, where numerous potential causal relations exist within these datasets.
We carefully select 1,438 videos (5.6k events) encompassing complicated causal relations and a wide range of scenarios as a new dataset: the MECD dataset, where 1,139 and 299 videos are randomly split for training and testing, respectively.
@@
Data Cleaning. As illustrated in ReXTime [18], only around ten percent of event pairs in daily video datasets such as ActivityNet Caption [45] are causal. Therefore, we further clean our dataset by excluding non-causal videos to guarantee the quality of our MECD dataset. We conducted data cleaning by our annotators with strict selection criteria: If more than 20% of the annotators believe that the video almost lacks causality, then the video will be excluded, e.g. video that describes multiple non-causal steps such as handcrafting.
@@
Dataset Annotation. The annotations of MECD dataset include 4 attributes. For videos sourced from the ActivityNet Captions dataset, The “duration”, “sentence”, and “timestamps” attributes in annotations remain the same as the original annotations for the dense video captioning task. For videos sourced from EgoSchema and NExTVideo, as lacking captioning and timestamps annotations, we adopt the approach of event-bench [47], dividing videos into 30-second segments as events and captioning each event using the state-of-the-art [48] video captioning model Gemini-Pro [49]. Additionally, our annotators further refine the “sentence” and “timestamps” attributes to improve the annotation quality.
@@
Specifically, in the context of our task, a new attribute, “causality”, is introduced. This attribute represents the causal relations between all premise events $\{e_{1},...,e_{N-1}\}$ and the final event $e_{N}$. To obtain this attribute, relations among events are first annotated using the GPT-4 API [50], and subsequently refined by five human annotators. Through a cross-annotation process [51, 52], ground truth causalities are determined by the majority of the annotators’ causal relation choices, thus mitigating potential inaccuracies and subjective biases to a certain extent. Annotation examples of MECD are shown in Fig. 4, our MECD dataset is carefully annotated to support the challenging task proposed with complete premise information.
@@
Additionally, to evaluate the model’s complete causal graph reasoning ability, an additional attribute, “complete causality”, is introduced for the test set, which represents all causal relations between any two events, and is annotated and refined in the same way as the “causality” attribute. Specifically, “complete causality” is a list of length (N-1), where N is the number of events in the video, and the k-th item (length = (k+1)) in this complete causality list represents the causal relations between $\{e_{1},...,e_{k+1}\}$ and $e_{k+2}$ as shown in Fig 4.
@@
We also present the ratio of causal relations of videos in Fig. 3 (b1), and the impact of events’ positions on their causality in Fig. 3 (b2). The proportion of events with or without causal relations is generally balanced, with the proportion of events without causal relations slightly higher. Moreover, by examining the last event as a case study, we investigated the correlation between causality and positional relations among events. Our findings indicate that the likelihood of adjacent events existing causal relations is higher, while the probability of other positions is approximately comparable.
@@ IV Method
In this section, we present our proposed Video Granger Causality Model (VGCM) to address Multi-Event Causal Discovery (MECD), as shown in Fig. 5. This model establishes the global connections across all events, and deduces the causality of a premise event by comparing the output features when it is masked or not, under the concept of the Event Causality Test. However, masking out an event may lead to the problem of confounding and illusion. In this context, we further utilize causal inference methods to address these by compensating or removing the effect of previous or subsequent causal events to mitigate the confounding meanwhile during inference the extra chain of thoughts and existence-only descriptions relieve the illusion. Additionally, we extend the modeling of event-causal relations to the global level and utilize context chain causal reasoning to increase the interaction of global information.

### v4

@@ I Introduction
Video causal reasoning aims to understand and analyze video content from a causal perspective. It requires models to comprehend temporal relationships, anticipate actions, and adapt to dynamic visual elements across frames. This capability is essential in various real-world applications, including autonomous driving [1], activity recognition [2], video surveillance [3], and even complex decision-making scenarios like robotic navigation [4]. Among these applications, Video Question Answering (VQA) [5, 6, 7] represents one of the most prominent tasks in video causal reasoning, where models are tested on their causal ability to understand videos through causal questions such as explanations, predictions, and counterfactual assumptions. Traditional VQA tasks can be viewed as attempts to discover single causal links within videos, as they typically identify one reason from given multiple options that explain the event described in the question.
Recent advancements in VQA have expanded beyond traditional tasks by incorporating more sophisticated causal reasoning methods. Notable trends include enhancing causal chain inference from simple one-to-one relations to multi-to-one chains, as seen in multiple correct answers VQA [8, 9, 10], and improving answer causal discovery through temporal grounding techniques, as seen in Grounded Video Question Answering (grounded VQA) [11, 12, 13]. Recent research also endeavors to infer a more comprehensive causal graph focusing on a particular event within its contexts [14, 15].
@@
Despite recent advancements, as shown in Tab. I, current video causal reasoning tasks remain limited in scope, primarily focusing on QA-based approaches that discover a single causal relation within a video. However, when the scene to be analyzed is more complex, these tasks often fail to meet the required understanding standards. For example, in a video where a person slips and falls while carrying a tray of drinks, the cause might be combined with a wet floor, slippery shoes, and losing balance due to walking too quickly. Besides, these tasks often lack the ability to perform fine-grained event-level reasoning, which is crucial in real-world scenarios. Such detailed reasoning typically leads to clearer insights into video content. Most importantly, they fail to provide a comprehensive, structured causal representation. For instance, in traffic surveillance videos, a detailed analysis of events occurring at different times is essential to determine which events, or combinations of events, resulted in a specific accident among the chain of traffic accidents.
@@
(1) Causality confounding indicates that the original causal relations between events are disrupted or interfered with by other relay or adjacent events. Such confounding stems from the fact that many causal relations flow through an intermediary event that acts as a bridge. For example, in Fig. 1(c), “submitting the paper” merely mediates the effect of “taking the test” on “obtaining a grade”, yet may be mistakenly treated as the sole cause, thereby obscuring the true causal dependency.
@@
(2) Illusory Causality, including both illusory temporal and existence causality. Illusory temporal causality exists when events exhibit a close correlation in temporal distribution. Such correlation may mislead the test of real causality. For example, in Fig. 1(d), “adding oil when cooking” often precedes “adding vegetables to stir-fry”, yet the former does not cause the latter. As for illusory existence causality, it occurs when some objects in early events may serve as necessary existence conditions of a later event. For instance, as shown in Fig. 1(e), while “a large brown dog enters the room” is a prerequisite for “the dog runs towards the camera”, the former does not cause the latter.
@@
Building upon the preceding discussion, we introduce a novel framework named VGCM (Video Granger Causality Model) to tackle MECD. This framework executes the Event Granger Test via an efficient mask-based event prediction model. According to the Granger Causality theorem in machine learning [17, 18, 19], if the inclusion of premise events can improve the prediction of a subsequent event, then they are causally related. Besides, the video data used in this article is theoretically and rigorously compliant with the conditions for Granger Causality, as it represents a continuous-in-time, deterministic, and complete system where potential confounders are mitigated through causal inference. Following this theory, VGCM deduces the causality of a premise event by comparing the predicted features of the result event when the premise is either masked or unmasked. Furthermore, to mitigate the challenges of causality confounding and illusory causality discussed earlier, we integrate two additional causal inference techniques—front-door adjustment [20, 21] and counterfactual inference [20, 22]—into our framework. Specifically, these techniques compensate for or remove the causal effects of previous or subsequent adjacent bridge events to eliminate confounding. Simultaneously, they address the illusory causality issue by incorporating the chain of thoughts [23, 24, 25] and existence-only descriptions.
Furthermore, we improve model robustness by applying context chain reasoning in the Event Causality Test, enhancing interactions among in-context causal chains. In reality, there are numerous instances where partial causes, when combined, contribute to the formation of the complete cause leading to the result. To enable the model to learn this reasoning pattern commonly found in reality, we randomly mask multiple premise events during training.
@@
This manuscript extends our NeurIPS 2024 paper [26] and makes the following contributions: The release and deeper analysis of MECD+, a larger multi-video source dataset. An enhancement of the Event Causality Test through robust context chain reasoning and efficient non-regressive complete graph reasoning. Additional metrics and further analysis of the MECD task, such as input modalities, hallucination issues, and downstream applications.
@@ II Related work
Video Causal Reasoning. Many existing tasks have tried to carry out causal reasoning in videos. Among these, the most common task is Video Question Answering (VQA), aiming to deduce a reason according to the question, grounded-QA methods such as SeViLA, NextGQA and Momentor [5, 11, 27] grounded a single reason happens before the result. However, the traditional VQA task does not extend to abducting multiple reasons, merely abducting a single causal chain from reason to the provided result. Some studies have enhanced video reasoning paths by introducing Chain-of-Thought (CoT) to better accomplish VideoQA [28, 29, 30] or Rumor Detection [31], however, these works remain constrained by single-chain causal inference. Recently, some methods, such as Glance [32], VideoTree [12], and TimeCraft [13] have constructed more detailed causal chains through reasoning from coarse-grained to fine-grained levels or bidirectional paths. Although these works improve the causal reasoning process, they only abduct the causal chain for a single result.
@@
Furthermore, many tasks are based on VQA for further causal reasoning attempts. Physical state reasoning task CLEVRER [7] and CATER [33] explored causal reasoning based on physics and other basic laws in virtual scenes. However, they haven’t been committed to extending to the general video. Neural-symbolic paradigm AAR [34] and LMLN [35] symbolized data and derived inference rules using external knowledge. However, they can only reason within a defined symbol domain. The most similar task VAR [15] predicted explanation events with premise events, and the causality is introduced during its prediction process. However, it hasn’t been committed to discovering the complete causal graph. Another concurrent study REXTIME [16] determined the causal relation between two events in a video and then designed QA questions targeted at the relations. However, it only reasoned through incomplete causal chains. Moreover, unlike the causal relation reasoning tasks mentioned above, the video spatial relationship reasoning task [36] aims at inferring the semantic connection between two objects formed by a specific action or state, rather than the causal relation that is our primary focus.
@@
The task of causal reasoning in natural language processing (NLP) generally involves inferring the causal relations between words within a text paragraph. TCR [43] proposes employing constrained conditional models and formulates the extraction of causal relations between events as an integer linear programming problem. The Following methods usually incorporate external knowledge, either through knowledge graphs or prompts. For example, RFBFN [44] and CauSeRL [45] utilized knowledge related to candidate causal events and external causal statements from the knowledge graph. However, these approaches are generally limited to handling sentences with a single causal relation pair. GESI [46] further advances this by constructing heterogeneous graph-based models to capture more complex causal relations, potentially spanning across multiple paragraphs. However, GESI suffers from slower inference speeds and increased complexity, which can be prohibitive for higher-dimensional video data applications. Additionally, following the Granger Causality method, CAUSE [47] proposes capturing event dependencies by fitting neural point processes and then employing axiomatic attribution to quantify the contribution of preceding events to subsequent predictions.
@@ III Multi-Event Causal Discovery
Figure 3: Statistics of the MECD Dataset. As shown in Figures (a1) and (a2), our dataset mainly analyzes videos that are longer than 120 seconds and contain five or more events. As shown in Figures (b1) and (b2), the proportion of causal and non-causal relations between events in the video of MECD is relatively balanced, moreover emphasizing the existence of causal relations between adjacent events.
To quantify the ability of causal discovery of a given model in multi-event videos, we propose the task of Multi-Event Causal Discovery (MECD). In formulation, given a video $\mathcal{E}$ that contains chronologically organized $N$ events, $\mathbb{E}:=\{e_{1},\dots,e_{N}\}$, this task aims at determining whether any previous event $e_{n}$ ($n<N$) has a causal relation with the last one (i.e., $e_{N}$). Specifically, an event $e_{n}=\{v_{n},c_{n}\}$ consists of a video clip $v_{n}$ and the corresponding caption $c_{n}$. Without loss of generality, relations of previous events to the last one can be expressed as ${\bm{r}}=[r_{1},\dots,r_{N-1}]$, where $r_{k}$ ($k<N$) is set to “1” to indicate the existence of $e_{k}$’s causal relation with $e_{N}$, and “0” otherwise. Notably, this setting (defined as causal chains discovery) is generalizable to causal relations of any two events as long as we cut off the video and treat the latter one as the last event.
@@ III-A Task Data
ActivityNet Captions dataset [48] is built on ActivityNet v1.3 which includes 20k YouTube videos. ActivityNet Captions is widely used in dense video captioning with annotated captions and timestamps of each sub-event. The EgoSchema dataset [49] contains over 5,000 video language understanding questions spanning over 250 hours of egocentric video data. NExTVideo [6] consists of 6,000 videos that are longer and richer in objects and interactions in daily life.
We carefully select 1,438 videos (5.6k events) encompassing complicated causal relations and a wide range of scenarios as a new dataset: the MECD dataset, where 1,139 and 299 videos are randomly split for training and testing, respectively. In our dataset, 1,106 videos are sourced from ActivityNet, 102 from EgoSchema, and 230 from NExTVideo.
@@
Data Cleaning. As illustrated in ReXTime [16], only around ten percent of event pairs in daily video datasets such as ActivityNet Caption [48] are causal. Therefore, we further clean our dataset by excluding non-causal videos to guarantee the quality of our MECD dataset. We conducted data cleaning by five annotators with strict selection criteria: If two or more annotators (out of five) label a video as containing at most one causal chain, then the video will be excluded, e.g. video that describes steps such as handcrafting. A large portion of videos lacking causal relations were excluded—43.5% from ActivityNet, 24.3% from NExTVideo, and 74.2% from EgoSchema.The results reveal that NExTVideo clips exhibit stronger causal coherence, while EgoSchema clips often lack it due to their ego-procedural nature.
@@
Dataset Annotation. The annotations of MECD dataset include 4 attributes. For videos sourced from the ActivityNet Captions dataset, The “duration”, “sentence”, and “timestamps” attributes in annotations remain the same as the original annotations for the dense video captioning task. For videos sourced from EgoSchema and NExTVideo, we select samples that are drawn from Event-Bench’s [50] collection, where all clips have been verified to contain discrete events within standardized 30-second intervals. We then caption each event using the SOTA [51] video captioning model Gemini-Pro [52]. Additionally, our annotators further refine the “sentence” and “timestamps” attributes to: (1) refine low-quality captions resulting from background clutter or ambiguous frames, and (2) improve event continuity that might be compromised by strict 30-second division.
@@
Specifically, in the context of our task, a new attribute, “causality”, is introduced. This attribute represents the causal relations between all premise events $\{e_{1},...,e_{N-1}\}$ and the final event $e_{N}$. To obtain this attribute, relations among events are first annotated using the GPT-4 API [53], and subsequently refined by five human annotators. Through a cross-annotation process [54, 55], ground truth causalities are determined by the majority of the annotators’ causal relation choices, thus mitigating potential inaccuracies and subjective biases to a certain extent. Besides, a Pearson correlation coefficient of 0.93 was achieved among the annotations, strongly indicating a high degree of objectivity in the annotation process. Our experiments also confirmed that the model’s ability to learn causality is robust to minor variations in the annotations. Annotation examples of MECD are shown in Fig. 4, our MECD dataset is carefully annotated to support the challenging task proposed with complete premise information.
@@
Furthermore, while the accuracy of deducing causality between the final event $e_{N}$ and its preceding events is a good general indicator, it doesn’t fully capture the model’s ability. To rigorously demonstrate the model’s capacity to deduce the complete causal graph, an additional attribute, “complete causality”, is introduced for the test set. This attribute represents all causal relations between any two events, and is annotated and refined in the same way as the “causality” attribute. Specifically, “complete causality” is a list of length (N-1), where N is the number of events in the video, and the k-th item (length = (k+1)) in this complete causality list represents the causal relations between $\{e_{1},...,e_{k+1}\}$ and $e_{k+2}$ as shown in Fig 4. The task of discovering the “complete causality” is defined as complete causal graph discovery.
@@
We also present the ratio of causal relations within every video in Fig. 3 (b1), and the impact of events’ positions on their causality in Fig. 3 (b2). The proportion of events with or without causal relations is generally balanced, with the proportion of events without causal relations slightly higher. Moreover, by examining the last event as a case study, we investigated the correlation between causality and positional relations among events. Our findings indicate that the likelihood of adjacent events existing causal relations is higher, while the probability of other positions is comparable.
@@ IV Method
In this section, we present our proposed Video Granger Causality Model (VGCM) to address Multi-Event Causal Discovery (MECD), as shown in Fig. 5. This model establishes the global connections across all events, and deduces the causality of a premise event by comparing the output features when it is masked or not, under the concept of the Event Causality Test. However, masking out an event may lead to the problem of confounding and illusion. We further employ causal inference methods to mitigate confounding by compensating for or removing the effects of preceding or subsequent causal events. Meanwhile, during inference, the extra chain of thoughts and existence-only descriptions help alleviate potential illusions. Additionally, we extend the modeling of event-causal relations to the global level and utilize context chain causal reasoning to increase the interaction of global information.